OpenStudy (anonymous):

use the FOIL method to solve: (x^2+7)(x^2+7)

OpenStudy (apoorvk):

@ParthKohli will FOIL it for you, he's an expert at this ;)

Parth (parthkohli):

lol

Parth (parthkohli):

FIRST \( \color{Black}{\Rightarrow x^2 \cdot x^2 }\) OUTSIDE \( \color{Black}{\Rightarrow x^2 \cdot 7 }\) INSIDE \( \color{Black}{\Rightarrow 7 \cdot x^2 }\) LAST \( \color{Black}{\Rightarrow 7 \cdot 7}\) Add them.

OpenStudy (anonymous):

i did all that already but i came up for my final answer with 16^2+49 is that right??

OpenStudy (anonymous):

please help!!

hero (hero):

FOIL METHOD and Distributive Method are the same exact thing

hero (hero):

(x^2+7)(x^2+7) =x^2(x^2+7)+7(x^2+7) =x^2 + 7x^2 + 7x^2 + 49 =x^2 + 14x^2 + 49

Parth (parthkohli):

Well... not really. They are the same thing, but they have different ways.

OpenStudy (anonymous):

but dont you need to collect like terms so itd be 15x^2+49

Parth (parthkohli):

And I find the way Hero did it easier.

hero (hero):

Parth, they're the same. Do you want me to prove it to you?

OpenStudy (anonymous):

dont argue answer my question please

Parth (parthkohli):

No lol they're same but done in different orders :P

OpenStudy (apoorvk):

They are the same actually, @Hero is right, just a 'tweak' different.

OpenStudy (apoorvk):

ya order of proceeding if a bit different.

OpenStudy (apoorvk):

*is

hero (hero):

look closely at the second step, then you'll realize they're exactly the same.

Parth (parthkohli):

The main rule is the same.

Parth (parthkohli):

Yeah Hero won the typing race ;)

OpenStudy (anonymous):

your answer was wrong hero.

Parth (parthkohli):

Yeah it must be x^4

hero (hero):

Oops

OpenStudy (anonymous):

thanks a lot dude -.-

hero (hero):

I blame Parth, lol

Parth (parthkohli):

lol

OpenStudy (anonymous):

makes me a sad panda

hero (hero):

Sorry, I wish I could make it up to you somehow

OpenStudy (anonymous):

(5x^3+5x)(x^2+6) answer this correctly and itll be okay xD

hero (hero):

(5x^3+5x)(x^2+6) =x^2(5x^3+5x)+6(5x^3+5x) =5x^5+5x^3+30x^3+30x =5x^5 + 35x^3 + 30x

OpenStudy (anonymous):

ok what about this one? (x^12-x^9)(x^3-10)

hero (hero):

Why haven't you figured out the FOIL Method yet?

OpenStudy (anonymous):

i have but im terrible with it and if i fail this quiz again for my online course im screwed. its just something i want to get done with

OpenStudy (anonymous):

is this right for it? x^15-11^12+10^9